explain the effect of temperature on rate of reaction?
 
 

Grade:12

2 Answers

pralay
8 years ago
The rate of reaction increases considerably with an increase in temperature. The rates of
many reactions are approximately doubled or tripled for every 10oC rise in temperature.
The temperature coefficient of a chemical reaction is defined as the ratio of the specific
reaction rates of a reaction at two temperature differing by 10oC.
μ = Temperature coefficient= k(r+10)/kt
Let temperature coefficient of a reaction be ' μ ' when temperature is raised from T1 to T2;
then the ratio of rate constants or rate may be calculated as
kT2/kT1 =(μ)(T2-T1)/10 = μ∆T/10
log10(kT2/kT1)= ∆T/10 log10μ
kT2/kT1 = antilog [∆T/10 log10μ]
Its value lies generally between 2 and 3.
 
SAI SARDAR
8 years ago
  1. TEMPERATURE IS DIRECTLY PROPORTIONAL TO THE RATE OF REACTION FOR NORMAL REACTIONS.
  2. IN EXOTHERMIC REACTIONS TEMPERATURE IS INCREASES RATE OF REACTION IS DECREASES.
  3. IN ENDOTHERMIC REACTION TEMPERATURE IS INCREASES RATE OF REACTION IS ALSO INCREASES.
              IF YOU LIKE MY ANSWER PLZ APPROVE IT.